what are the intercepts of the equation 7x-2y-14z=56
Answer by Fombitz(32388)

x-intercept, +y=z=0
+7x-2y-14z=56
+7x=56
+x=8
(8,0,0)
.
.
.
y-intercept, x=z=0
++7x-2y-14z=56
+-2y=56
+y=-28
(0,-28,0)
.
.
.
z-intercept, +x=y=0
++7x-2y-14z=56
+-14z=56
+z=-4
(0,0,-4)